OpenStudy (station22):

Determine the prime power representations of both 165 and 297

OpenStudy (zzr0ck3r):

\(165=5*33=5*11*3\)

OpenStudy (zzr0ck3r):

do something similar for the other one.

OpenStudy (sshayer):